# Universal scaling law for quantum droplet formation

### v1

@@ Abstract
Given the right set of circumstances, ultracold quantum gases are able to change character and condense into a liquid state of quantum droplets. The size distribution of the droplets is determined dynamically in the condensation process. A semi-quantitative argument is presented which suggests that, at zero temperature, a multiple droplet system has is a preferred scale $\propto v^{-d}$, with $d\approx 0.5$, where $v$ is the rate of change of parameters at the time of droplet formation. Numerical simulations of two dimensional systems strongly support a power law, but with an exponent $d\in(0.3,0.4)$.
@@
The scaling law for droplet sizes resembles the Kibble-Zurek scaling law for structure formation in second order phase transitions Kibble (1976); Zurek (1985); del Campo and Zurek (2014), which involves a power of the rate parameter. There is also good evidence that a Kibble-Zurek type of scaling law applies to some first order phase transitions in cold atom systems Panagopoulos and Vicari (2015); Kang et al. (2017); Liang and Zhong (2017); Coulamy et al. (2017); Qiu et al. (2020). The physical origin of the scaling laws for the droplets has some similarities, in that there is a limitation on the size of structures that can form on some timescale. In the thermal case, this is set by the relaxation time of the medium. In the case of droplet formation, which can occur at zero temperature in the absence of dissipation, this is related to the growth rate of the droplet instability.
@@
Thus, unstable modes exist when $n_{d}>\bar{n}$ and $V^{\prime\prime}_{LHY}$ becomes negative. Since these instabilities mark the onset of droplet formation, it seems reasonable to suppose their wavelengths will be related to the distance between droplets. The growth rate of these modes, $v_{d}={\rm Im}(\omega)$ is a function of wave number and time. It has a maximum at a given time for a particular wave number $k_{\rm max}$,
$v_{d}=\frac{\hbar k_{\rm max}^{2}}{2m}=\frac{1}{\hbar}|\bar{n}V_{LHY}^{\prime\prime}|.$ (3)
Now consider what happens to a uniform system that starts in the stable regime with $n_{d}<\bar{n}$, following which the physical parameters are slowly changed so that $n_{d}$ increases at a rate $v$. Initially, the growth rate $v_{d}=0$, and there is no droplet formation. When $n_{d}\geq\bar{n}$, the growth rate $v_{d}$ slowly increases from zero. The growth rate becomes comparable to the system timescale when $v_{d}\sim v$. Taking $v_{d}=v$ sets a critical length scale $2\pi/k_{\rm max}$ from Eq. (3), which we identify as the preferred separation between neighbouring droplets $r_{\rm sep}$. The conclusion is that the droplet separations scale with the rate $v$ as
$r_{\rm sep}\propto v^{-1/2}.$ (4)
Knowing the mean density of the system $\bar{n}$, the density of a typical droplet $n_{d}$ and the separation between droplets fixes the droplet radius $r_{d}$, $r_{d}=r_{\rm sep}\sqrt{\bar{n}/\pi n_{d}}\propto v^{-d}$. There is a caveat for large rates $v$, when the right hand term in Eq. (3) can have a maximum value $v_{\rm max}<v$, in which case the droplet separations would be set by $v_{\rm max}$ and would become independent of $v$.
@@
The relation between droplet sizes and parameter evolution rates is shown in figure 2. The values of the parameters $g^{\prime}/g$ and $n_{0}\xi_{0}^{2}$ differ by factors up to 4 between the plots to show that the scaling relation goes beyond a trivial rescaling invariance of the GPE. The data are excellent fits to power-law relations $r_{d}\propto v^{-d}$. The small variation in the value of $d$ for large changes in regime is consistent with small effects due to droplet mergers. However, the power $d\in(0.3,0.4)$ is significantly smaller than the naive prediction $d=0.5$ given earlier.
@@
Conclusions: Ultra cold quantum gases condense into liquid quantum droplets under certain parameter regimes. In a many droplet system, the size distribution of these droplets is determined dynamically in the droplet transition. It has been argued here that there is a refered droplet size and separation fixed in terms of the rate of change of parameters $v$ by $r_{d}\propto v^{-d}$, where $d\approx 1/2$. This is independent of the detailed nature of the system.
Numerical simulations show strong evidence for a power law, but with $d\in(0.3,0.4)$. It would be interesting to pursue this further beyond two dimensions, to examine the effects of thermal fluctuations, and to consider “unbalanced” mixtures Flynn et al. (2023). The proposed range of experimental parameters look challenging, but further numerical analysis may allow some movement in these values
@@
Data supporting this publication are openly available under a Creative Commons CC-BY-4.0 Licence in Dat (2025)
@@ Acknowledgements.
The author is grateful for helpful comments from Tom Billam, Matt Johnson, Markus Oberthaler and Yansheng Zhang. This work is supported by the UK Science and Technology Facilities Council [grants ST/T00584X/1 and ST/W006162/1]. The author is also grateful for the hospitality of Perimeter Institute where part of this work was carried out. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities.
Appendix A Supplementary Material

### v2

@@ Abstract
Given the right set of circumstances, ultracold quantum gases are able to change character and condense into a liquid state of quantum droplets. The size distribution of the droplets is determined dynamically in the condensation process. A semi-quantitative argument is presented which suggests that, at zero temperature, a multiple droplet system has is a preferred scale $\propto v^{-1/3}$, where $v$ is the rate of change of parameters at the time of droplet formation. Numerical simulations of two dimensional systems strongly support a power law $v^{-d}$, with an exponent $d\in(0.327,0.375)$.
@@
The scaling law for droplet sizes resembles the Kibble-Zurek scaling law for structure formation in second order phase transitions Kibble (1976); Zurek (1985); del Campo and Zurek (2014), which involves a power of the rate parameter. The physical origin of the scaling laws for the droplets has some similarities, in that there is a limitation on the size of structures that can form on some freeze-out timescale. In the thermal case, this is set by the relaxation time of the medium. In the case of droplet formation, which can occur in the absence of dissipation, the process is similar to the “underdamped” Kibble-Zurek effect Suzuki and Zurek (2024). However, an important difference is that droplet formation can happen at zero temperature from purely quantum fluctuations that cannot be modelled by a Langevin equation approach. There is also good evidence that a Kibble-Zurek type of scaling law applies to other situations, for example some first order phase transitions in cold atom systems Panagopoulos and Vicari (2015); Kang et al. (2017); Liang and Zhong (2017); Coulamy et al. (2017); Qiu et al. (2020).
@@
Thus, unstable modes exist when $n_{d}>\bar{n}$ and $V^{\prime\prime}_{LHY}$ becomes negative, marking the onset of droplet formation. If the parameters change at a rate $v$ near the droplet transition, then we can replace $V^{\prime\prime}_{LHY}$ by $-gvt$, where $g$ is related to the third derivative of $V_{LHY}$ and the instability sets in at $t=0$.
Following Sazuki and Zurek Suzuki and Zurek (2024), finding the characteristic length scales can be broken down into two steps. The first step is to find the freeze-out time when the instability becomes non-linear, and the second step is to convert this into a length scale. Taking the dispersion relation (2) as a starting point, the Bogoliubov modes will satisfy a non-linear equation of the form
$m\delta\ddot{\psi}-gnk^{2}vt\delta\psi+O(\delta\psi^{2})=0,$ (3)
in the small $k$ limit. The $v$ dependence can be removed by introducing a time coordinate $\tilde{t}=v^{1/3}t$, and rescaling $\delta\psi$ to remove any $v$ dependence in the $\delta\psi^{2}$ terms. The solutions for any given $k$ mode are independent of $v$ and become non-linear at some fixed value for the rescaled time $\tilde{t}(k)$.
Returning to the dispersion relation (2), there is a value of $k=k_{\rm max}$ for which the growth rate $\Im(\omega)$ reaches a maximum [^2] The freeze-out time was found using a long wavelength approximation which is relaxed when finding the length scale. This is consistent with the approach used in Ref. Suzuki and Zurek (2024), $k_{\rm max}\propto|V_{LHY}^{\prime\prime}|^{1/2}\propto(vt)^{1/2}$. We may interpret the corresponding length scale as the droplet separation $r_{s}$, $r_{s}\propto(vt)^{-1/2}$. Since the freeze-out time $t=v^{-1/3}\tilde{t}$, it follows that
$r_{s}\propto v^{-1/3}.$ (4)
Knowing the mean density of the system $\bar{n}$, the density of a typical droplet $n_{d}$ and the separation between droplets fixes the droplet radius $r_{d}$, $r_{d}=r_{\rm sep}\sqrt{\bar{n}/\pi n_{d}}\propto v^{-1/3}$.
@@
The relation between droplet sizes and parameter evolution rates is shown in figure 2. The values of the parameters $g^{\prime}/g$ and $n_{0}\xi_{0}^{2}$ differ by factors up to 4 between the plots to show that the scaling relation goes beyond a trivial rescaling invariance of the GPE. The data are excellent fits to power-law relations $r_{d}\propto v^{-d}$. The small variation in the value of $d$ for large changes in regime is consistent with small effects due to droplet mergers. However, the power $d\in(0.327,0.375)$ is consistent with the naive prediction $d=1/3$ given earlier.
@@
Conclusions: Ultra cold quantum gases condense into liquid quantum droplets under certain parameter regimes. In a many droplet system, the size distribution of these droplets is determined dynamically in the droplet transition. It has been argued here that there is a refered droplet size and separation fixed in terms of the rate of change of parameters $v$ by $r_{d}\propto v^{-d}$, where $d=1/3$. This is independent of the detailed nature of the system.
Numerical simulations show strong evidence for a power law, with $d\in(0.327,0.375)$. It would be interesting to pursue this further beyond two dimensions, to examine the effects of thermal fluctuations, and to consider “unbalanced” mixtures Flynn et al. (2023). The proposed range of experimental parameters look challenging, but further numerical analysis may allow some movement in these values
@@
Data supporting this publication are openly available under a Creative Commons CC-BY-4.0 Licence in Dat
@@ Acknowledgements.
The author is grateful for helpful comments from Tom Billam, Matt Johnson, Markus Oberthaler, Yansheng Zhang and Mark Hindmarsh. This work is supported by the UK Science and Technology Facilities Council [grants ST/T00584X/1 and ST/W006162/1]. The author is also grateful for hospitality of the Perimeter Institute and the University of Helsinki where part of this work was carried out. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities.
Supplementary Material